# Reliable Fine-Grained Evaluation of Natural Language Math Proofs

### v1

@@ Abstract
Recent advances in large language models (LLMs) for mathematical reasoning have largely focused on tasks with easily verifiable final answers; however, generating and verifying natural language math proofs remains an open challenge. We identify the absence of a reliable, fine-grained evaluator for LLM-generated math proofs as a critical gap. To address this, we propose a systematic methodology for developing and validating evaluators that assign fine-grained scores on a 0–7 scale to model-generated math proofs. To enable this study, we introduce ProofBench, the first expert-annotated dataset of fine-grained proof ratings, spanning 145 problems from six major math competitions (USAMO, IMO, Putnam, etc) and 435 LLM-generated solutions from Gemini-2.5-pro, o3, and DeepSeek-R1. Using ProofBench as a testbed, we systematically explore the evaluator design space across key axes: the backbone model, input context, instructions and evaluation workflow. Our analysis delivers ProofGrader, an evaluator that combines a strong reasoning backbone LM, rich context from reference solutions and marking schemes, and a simple ensembling method; it achieves a low Mean Absolute Error (MAE) of 0.926 against expert scores, significantly outperforming naive baselines. Finally, we demonstrate its practical utility in a best-of-$n$ selection task: at $n=16$, ProofGrader achieves an average score of 4.14/7, closing 78% of the gap between a naive binary evaluator (2.48) and the human oracle (4.62), highlighting its potential to advance downstream proof generation.
@@ 1 Introduction
We identify the absence of a reliable proof evaluator as a key bottleneck for improving proof generation, which is essential for providing faithful assessments of model capabilities and accurate reward signals for training models. Expert grading, while accurate, is slow and costly. While formal math (e.g., Lean) offers absolute certainty, it remains detached from the natural language used in most human mathematics education and research; furthermore, automatically translating natural-language proofs into formal languages is brittle and remains extremely challenging (Gao et al., 2025; Liu et al., 2025a). Our work therefore focuses on the critical and complementary task of evaluating proofs in their natural representation. While LLM-as-a-judge (Zheng et al., 2023; Sheng et al., 2025; Dekoninck et al., 2025; Huang & Yang, 2025) is promising, its application to math proofs is unsettled, and outcomes are sensitive to evaluator design—model choice, available context, rubric construction, and prompting—all of which are poorly understood.
@@
Finally, we validate the evaluators’ practical utility in a downstream best-of-$n$ selection task, a standard proxy for assessing its potential as a reward signal (Gao et al., 2022; Frick et al., 2024; Malik et al., 2025). At $n=16$, ProofGrader achieves an average score of 4.14/7, closing 78% of the performance gap between a naive binary evaluator (2.48) and the human oracle (4.62). It also outperforms computationally intensive, pairwise selection methods such as Knockout tournament selection (Liu et al., 2025b). These results highlight ProofGrader’s promise as a reward model for advancing proof generation.
To summarize, our contributions are three-fold:
  • We introduce ProofBench [^1] The full dataset is available at huggingface.co/datasets/wenjiema02/ProofBench., an expert-graded math proof dataset consisting of problems from well-established contests (USAMO, IMO, Putnam, etc) and solutions from state-of-the-art reasoning models (o3, gemini-2.5-pro and DeepSeek-R1) (§2).
@@ 2 ProofBench: Expert-Rated Math Proof Solutions
We assess proof evaluators for human alignment on a 0–7 scale, which requires fine-grained expert annotations on model-generated proofs across diverse sources and years. Prior work such as MathArena focuses on a small set of contests like USAMO 2025 (Petrov et al., 2025); other studies of evaluators are limited to binary judgments of proof correctness (Dekoninck et al., 2025). We thus construct our own dataset, ProofBench, consisting of 145 proof problems from six major math competitions across four years, with model-generated solutions by the state-of-the-art reasoning models: o3, Gemini-2.5-Pro, and DeepSeek-R1. This section describes the data collection, annotation process (§2.1), and dataset statistics (§2.2).
@@ 2.1 Dataset Construction
We collected 145 problems from the official websites of prestigious mathematics competitions, including the APMO, EGMO, IMO, PUTNAM, USA TST, and USAMO, spanning the years 2022–2025. Distribution of contests is shown in Figure 3a. Sourcing directly from official materials ensures the reliability of problem statements and ground-truth solutions, avoiding transcription errors common in secondary sources. Problems were parsed from official PDFs and normalized, and all available human solutions were included. The final collected data consists of a set of (problem, reference solution(s), metadata) triples, where metadata includes the source competition and year.
@@
For each problem, we generate a proof from models (generators) using a standardized prompt (§A.7) that asks for a complete, self-contained proof. We generate proofs from three state-of-the-art reasoning models: OpenAI o3, Gemini-2.5-Pro, and DeepSeek-R1-0528, which span both proprietary and open-source families.
Expert Annotations.
@@
Step 1: Automated Marking Scheme Generation: A central challenge in creating this benchmark was ensuring consistent and scalable expert grading. While human-written rubrics are the gold standard, automated generation offers superior scalability and a systematic way to recognize key steps from provided reference solutions. In our method, the marking scheme is produced by an LLM $\mathcal{M}_{\text{MS}}$. Given the problem $x$ and reference solutions $\mathcal{S}$, $\mathcal{M}_{\text{MS}}$ is prompted to output (i) a list of conditions under which scores are awarded or deducted, and (ii) a list of trivial cases that should not be awarded any points. This generator was developed through a rigorous refinement process where experts also graded the quality of the generated marking schemes, providing feedback that led to the selection of gemini-2.5-pro as the final $\mathcal{M}_{\text{MS}}$ (see §A.2 for details). An example of a generated marking scheme together with problem text and reference solution is shown in Figure 2. From our expert evaluations, approximately 85% of the generated marking schemes in our final dataset were judged to be reasonable and of high quality.
Step 2: Proof Grading: With a problem-specific marking scheme, an expert annotator scores a given model-generated proof. Experts were instructed to treat the marking scheme as a detailed reference for the expected solution path, rather than a rigid checklist. This is particularly important for fairly evaluating proofs that employed a novel method different from the provided ground-truth solutions, ensuring that valid alternative reasoning paths were credited appropriately. The experts will assign a score on a 0–7 scale. Before large-scale annotation, our experts underwent a calibration phase to align their judgments and finalize rules for handling edge cases, a critical step for minimizing inconsistencies. We assign 40% of the solutions to more than one annotator and ensure their agreement is 90% or higher.
Figure 3: Data statistics and model evaluation results. (a) Problem statistics across six competitions. (b) Score distribution showing mean (2.41) and median (1.00). (c) Model performance comparison via box plots. Each box represents the interquartile range (IQR, middle 50% of scores), with the red line showing the median and blue dashed line showing the mean. (d) Per-generator score distribution. Stacked horizontal bar chart showing the percentage of problems receiving each score (0-7) for each model. Colors indicate score values (red=low, green=high). (e) Performance heatmap by generator and competition. (f) Competition difficulty ranking with error bars: TST is the most challenging competition (mean: 1.28) and Putnam is the easiest (mean: 3.11).
@@ 2.2 Dataset Statistics
In total, ProofBench consists of 145 proof problems and 435 expert-annotated evaluations of proofs generated by OpenAI o3, Gemini-2.5-Pro, and DeepSeek-R1-0528. Figure 3 summarizes data statistics across competitions (Figure 3a), scores (Figure 3b), and generators (Figure 3c).
@@
  • Current models remain far from achieving high scores: even the strongest models obtain scores of $6$ or higher on fewer than 30% of problems, as shown in Figure 3d.
  • OpenAI-o3 performs best overall, with a mean score of 2.87, and is closely followed by Gemini-2.5-Pro, according to Figure 3c.
  • Performance varies considerably across competitions (Figure 3e,f): all models perform best on Putnam, where the mean score is 3.11, but struggle on TST, where the mean drops to 1.28.
@@ 3 A Systematic Study of Evaluator Designs
Using ProofBench, we systematically study the key design factors for a math proof evaluator that produces fine-grained scores (0–7) for model-generated solutions. We demonstrate that (i) a strong reasoning backbone with informative context (like marking schemes) yields substantial gains, (ii) simple ensembling further improves accuracy and robustness, and (iii) a staged evaluation pipeline can refine weaker model backbones.
@@ 3.1 Evaluator Designs
Given the task of scoring model-generated solutions, a naive evaluator would simply take a problem and a generic instruction and assign a score. Instead, we systematically investigate design choices that enhance evaluation quality beyond this baseline. We first analyze single-pass evaluators along four key dimensions: (i) backbone model, (ii) contextual input, (iii) instruction set, and (iv) workflow design. We then extend this design to ensemble-based and multi-stage evaluation workflows.
@@
A single-pass evaluator prompts a backbone model $\mathcal{M}$ (which may or may not be the same as $\mathcal{M}_{\text{MS}}$) to grade a solution $s$ in one step. We analyze single-pass evaluator along three dimensions: the backbone LM, the context provided and the instructions given.
  • Backbone Model Choice. This refers to the LLM that is prompted to execute the evaluation. We compare six models that span different model families, size and reasoning capabilities: o3, gpt-5 (GPT-5-Thinking), Gemini (gemini-2.5-pro), o4-mini, R1 (DeepSeek-R1-0528), and gpt-4o.
@@
  • Instruction. The instruction set refers to the specific prompt that guides the evaluator on how to interpret and apply the provided context information and perform the task. In our study, for the most informative context setting, Ref+MS, we compared three types of instructions to understand how guidance on using the context affects performance. These include Norm (Normal), a flexible instruction that directs the model to follow the marking scheme but allows it to map valid alternative approaches to equivalent checkpoints; Strict, a rigid instruction that requires the model to adhere strictly to the provided marking scheme and penalize any deviation; and Basic, an instruction with minimal guidance on how to use the provided materials.
We then further study whether extending single-pass evaluators, e.g., through ensembling or multi-stage evaluation workflows, improves performance.
@@
Table 1: Performance comparison of six LLMs on 0-7 scale evaluation tasks under different context designs. Contexts: Ref+MS (reference solution + marking scheme), MS (marking scheme only), Ref (reference solution only), None (no context). Values shown as mean $\pm$ 95% confidence interval (CI) margin of error. Best values per model in bold; best ($\blacksquare$) and worst ($\blacksquare$) context highlighted. Arrows: $\downarrow$ lower better, $\uparrow$ higher better, $\approx$ closer to zero better. Quality: $\blacksquare$ Excellent, $\blacksquare$ Good, $\blacksquare$ Fair, $\blacksquare$ Poor, $\blacksquare$ Bad.
@@ 3.2 Evaluation Metrics
MAE and RMSE both measure errors relative to expert scores; MAE averages absolute deviations, while RMSE takes the square root of mean squared deviations and therefore penalizes large mistakes more (lower is better). WTA(1)($\leq$ 1) measures the fraction of predictions that land within one point of the expert score. Bias measures the average signed error (systematic shift), with positive values indicating over-scoring and negative values indicating under-scoring.
For ranking agreement within a problem, we use Kendall’s $\tau_{b}$ (ties-adjusted). Detailed definitions of Kendall’s $\tau_{b}$ as well as aggregation formulas are provided in §A.3.
@@ 3.3 What Factors Improve Single-pass Evaluator?
Evaluator MAE $\downarrow$ by Generator Outputs
R1 Gemini o3
o3 0.993 0.778 1.120
Gemini $1.183$ $1.616$ $1.225$
R1 $1.504$ $1.432$ $1.181$
Table 3: Per-generator evaluation accuracy (MAE; lower is better). Bold marks, for each generator (column), the best (smallest) MAE across generators.
@@
A natural question that can arise is the potential bias when evaluators assess outputs generated by their own models: are they more accurate because of familiarity, or less so because they overestimate their own generations? To investigate this, we evaluate three single-pass evaluators, each using the same reference solutions and marking scheme, on per-generator splits, i.e., response sets grouped by their source generator (R1, Gemini, o3). The results in Table 3 are mixed. On one hand, the strongest evaluator remains consistent across all generators: the o3-based evaluator performs best across all of them. On the other hand, each evaluator exhibits its highest MAE on outputs produced by its own model family, indicating a tendency toward within-generator underperformance. Per-generator significance tests and additional breakdowns are provided in §A.4.
@@ 3.4 Do Ensembling and Staged Workflows Improve Single-pass Evaluation?
We next extend single-pass evaluators through either ensembling or multi-stage pipelines, as described in §3.1.
@@
Although the absolute performance gains may appear modest, the key advantage of ensembling is variance reduction: single-pass evaluators, even with rich context, can exhibit high variance, whereas ensembling substantially mitigates it.
@@
o3 Single-pass (Ref+MS) $1.273$ $0.964$ $76.5$ $0.502$ 0.008-0.008
Binary+Errors $\rightarrow$ Fine-Grained $1.375$ $1.065$ $73.0$ $0.444$ 0.102-0.102
o4-mini Single-pass (Ref+MS) $1.816$ $1.367$ $67.6$ $0.476$ $0.762$
Binary+Errors $\rightarrow$ Fine-Grained $1.650$ $1.245$ $66.7$ $0.503$ $0.320$
o3+o4-mini Binary+Errors $\rightarrow$ Fine-Grained $1.513$ $1.126$ $71.4$ $0.458$ 0.095-0.095

### v2

@@ Abstract
Recent advances in large language models (LLMs) for mathematical reasoning have largely focused on tasks with easily verifiable final answers while generating and verifying natural language math proofs remain an open challenge. We identify the absence of a reliable, fine-grained evaluator for LLM-generated math proofs as a critical gap. To address this, we propose a systematic methodology for developing and validating evaluators that assign fine-grained scores on a 0–7 scale to model-generated math proofs. To enable this study, we introduce ProofBench, the first expert-annotated dataset of fine-grained proof ratings, spanning 145 problems from six major math competitions (USAMO, IMO, Putnam, etc.) and 435 LLM-generated solutions from Gemini-2.5-Pro, o3, and DeepSeek-R1. Using ProofBench as a testbed, we systematically explore the evaluator design space across key axes: the backbone model, input context, instructions and evaluation workflow. Our analysis delivers ProofGrader, an evaluator that combines a strong reasoning backbone LM, rich context from reference solutions and marking schemes, and a simple ensembling method; it achieves a low Mean Absolute Error (MAE) of 0.926 against expert scores, significantly outperforming naive baselines. Finally, we demonstrate its practical utility in a best-of-$n$ selection task: at $n=16$, ProofGrader achieves an average score of 4.14/7, closing 78% of the gap between a naive binary evaluator (2.48) and the human oracle (4.62), highlighting its potential to advance downstream proof generation.
@@ 1 Introduction
We identify the absence of a reliable proof evaluator as a key bottleneck for improving proof generation, which is essential for providing faithful assessments of model capabilities and accurate reward signals for training models. Expert grading, while accurate, is slow and costly. While formal math (e.g., Lean) offers absolute certainty, it remains detached from the natural language used in most human mathematics education and research; furthermore, automatically translating natural-language proofs into formal languages is brittle and remains extremely challenging (Gao et al., 2025; Liu et al., 2025a). Our work therefore focuses on the critical and complementary task of evaluating proofs in their natural representation. While LLM-as-a-judge (Zheng et al., 2023b; Sheng et al., 2025; Dekoninck et al., 2025; Huang and Yang, 2025) is promising, its application to math proofs is unsettled, and outcomes are sensitive to evaluator design—model choice, available context, rubric construction, and prompting—all of which are poorly understood.
@@
Finally, we validate the evaluators’ practical utility in a downstream best-of-$n$ selection task, a standard proxy for assessing its potential as a reward signal (Gao et al., 2022; Frick et al., 2024; Malik et al., 2025). At $n=16$, ProofGrader achieves an average score of 4.14/7, closing 78% of the performance gap between a naive binary evaluator (2.48) and the human oracle (4.62). It also outperforms computationally intensive, pairwise selection methods such as Knockout tournament selection (Liu et al., 2025b). These results highlight ProofGrader’s effectiveness in identifying high-quality proofs. To summarize, our contributions are three-fold:
  • We introduce ProofBench [^1] The full dataset is available at huggingface.co/datasets/wenjiema02/ProofBench., an expert-graded math proof dataset consisting of problems from well-established contests and solutions from state-of-the-art reasoning models (§2).
@@ 2 ProofBench: Expert-Rated Math Proof Solutions
We assess proof evaluators for human alignment on a 0–7 scale, which requires fine-grained expert annotations on model-generated proofs across diverse sources and years. Existing resources are limited in scale or scoring granularity (Petrov et al., 2025; Dekoninck et al., 2025). We therefore construct our own dataset, ProofBench, consisting of 145 proof problems from six major math competitions across four years, with model-generated solutions by the state-of-the-art reasoning models: o3, Gemini-2.5-Pro, and DeepSeek-R1-0528. This section describes the data collection, annotation process, and statistics; §3 then describes how we use it to evaluate proof evaluators.
@@ 2.1 Dataset Construction
We collected 145 problems from the official websites of prestigious mathematics competitions, including the APMO, EGMO, IMO, Putnam, USA TST, and USAMO, spanning the years 2022–2025. Whenever possible, problems were parsed from official PDFs, normalized, and accompanied by all available human-written solutions. This ensures the reliability of both problem statements and ground-truth solutions, avoiding transcription errors commonly found in secondary sources. When official solutions were unavailable, we supplemented the data with materials from reliable third-party websites. More details about each competition are provided in §B.
(a) Problem Distribution
(b) Overall Score Distribution
(c) Competition Difficulty
(d) Model vs. Competition
(e) Per-model Score Breakdown
(f) Contamination Analysis
Figure 3: Dataset information. (a) Distribution of unique problems across the six competitions. (b) Aggregate histogram of scores (0–7) across all models. (c) Competition difficulty ranking based on average scores; Putnam yielded the highest performance (3.09), while TST was the most challenging (1.26). (d) Heatmap of average scores by model and competition (Green indicates higher performance). (e) Stacked bar chart detailing the percentage of problems achieving each score (Red=0, Green=7). (f) Per-model box plots comparing performance on problems released before vs. after the model’s knowledge cutoff. There is a consistent drop in mean and median scores post-cutoff.
@@
For each problem, we generate a proof from models (generators) using a standardized prompt (§K) that asks for a complete, self-contained proof. We generate proofs from three state-of-the-art reasoning models: OpenAI o3 (OpenAI, 2025b), Gemini-2.5-Pro (Google DeepMind, 2025), and DeepSeek-R1-0528 (DeepSeek-AI et al., 2025), which span both proprietary and open-source families.
Expert Grading Pipeline.
@@
Stage 1: Marking Scheme Finalization. A central challenge in creating this benchmark was ensuring consistent and scalable expert grading. While human-written rubrics are the gold standard, automated generation offers superior scalability and a systematic way to recognize key steps from provided reference solutions. In our method, the marking scheme is produced by an LLM $\mathcal{M}_{\text{MS}}$. Given the problem $x$ and reference solutions $\mathcal{S}$, $\mathcal{M}_{\text{MS}}$ is prompted to output (i) a list of conditions under which scores are awarded or deducted, and (ii) a list of trivial cases that should not be awarded any points. This generator was developed through a rigorous refinement process where experts also graded the quality of the generated marking schemes, providing feedback that led to the selection of gemini-2.5-pro as the final $\mathcal{M}_{\text{MS}}$ (see §C for details). An example of a generated marking scheme together with problem text and reference solution is shown in Figure 2. From our expert evaluations, approximately 85% of the generated marking schemes in our final dataset were judged to be reasonable and of high quality.
Stage 2: Proof Grading. With a problem-specific marking scheme, an expert annotator scores a given model-generated proof. Experts were instructed to treat the marking scheme as a detailed reference for the expected solution path, rather than a rigid checklist. This is particularly important for fairly evaluating proofs that employed a novel method different from the provided ground-truth solutions, ensuring that valid alternative reasoning paths were credited appropriately. The experts will assign a score on a 0–7 scale. Before large-scale annotation, our experts underwent a calibration phase to align their judgments and finalize rules for handling edge cases, a critical step for minimizing inconsistencies. We assign 41% of the solutions to more than one annotator and their within-1-point agreement is 87.5%. Disagreements are further resolved through discussions.
@@ 2.2 Dataset Statistics
In total, ProofBench consists of 145 proof problems and 435 expert-annotated evaluations of proofs generated by OpenAI o3, Gemini-2.5-Pro, and DeepSeek-R1-0528. Figure 3 summarizes the dataset statistics, including the problem distribution across competitions (3(a)) and the aggregate score frequencies (3(b)).
@@
  • As shown in 3(e), current models remain far from reliable; even the strongest models achieve scores of 66 or higher on fewer than 30% of problems. OpenAI o3 demonstrates the strongest overall performance with the highest mean scores in 5/6 competitions (3(d)).
  • Performance varies considerably across problem sources (3(c)3(d)). Models perform best on Putnam (mean 3.09) but struggle markedly on TST (mean 1.26).
  • Comparison of performance on problems released before versus after the knowledge cutoff (3(f)) reveals a consistent drop in scores on newer problems, suggesting potential contamination in the training data or a generalization gap on novel problems.
@@ 3 A Systematic Study of Evaluator Designs
Using ProofBench, we systematically study the key design factors for a math proof evaluator that produces fine-grained scores (0–7) for model-generated solutions. We demonstrate that (i) a strong reasoning backbone with informative context (like marking schemes) yields substantial gains, and (ii) simple ensembling further improves accuracy and robustness.
@@ 3.1 Evaluator Designs
Given the task of scoring model-generated solutions, a naive evaluator would simply take a problem and a generic instruction and assign a score. Instead, we systematically investigate design choices that enhance evaluation quality beyond this baseline. We first analyze single-pass evaluators along three key dimensions: (i) backbone model, (ii) contextual input, and (iii) instruction set. We then extend this design to ensemble-based and multi-stage evaluation workflows.
@@
A single-pass evaluator prompts a backbone model $\mathcal{M}$ (which may or may not be the same as $\mathcal{M}_{\text{MS}}$) to grade a solution $s$ in one step. We analyze single-pass evaluators along three dimensions: the backbone LM, the context provided and the instructions given.
  • Backbone Model Choice. This refers to the LLM that is prompted to execute the evaluation. We compare six models that span different model families, sizes and reasoning capabilities: o3, gpt-5 (GPT-5-Thinking (OpenAI, 2025a)), Gemini (Gemini-2.5-Pro), o4-mini (OpenAI, 2025b), R1 (DeepSeek-R1-0528), and gpt-4o (OpenAI, 2024).
@@
  • Instruction. The instruction set refers to the prompt that guides the evaluator on how to interpret and apply the provided context and perform the task. In our study, for the most informative context setting, Ref+MS, we compared 3 types of instructions to understand how guidance on using the context affects performance. These include Norm (Normal), a flexible instruction that directs the model to follow the marking scheme but allows it to map valid alternative approaches to equal checkpoints; Strict requires the model to adhere strictly to the provided marking scheme and penalize any deviation; and Basic has minimal guidance on how to use the provided materials.
@@
Table 1: Performance comparison of 6 LLMs on 0-7 scale evaluation tasks under different context designs. Contexts: Ref+MS (reference + marking scheme), MS (marking scheme only), Ref (reference only), None (no context). Values shown as mean with $\pm$ 95% CI. Best values per model in bold; best ($\blacksquare$) and worst ($\blacksquare$) context highlighted. Arrows: $\downarrow$ lower better, $\uparrow$ higher better, $\approx$ closer to zero better. Quality: $\blacksquare$ Excellent, $\blacksquare$ Good, $\blacksquare$ Fair, $\blacksquare$ Poor, $\blacksquare$ Bad.
@@ 3.2 Evaluation Metrics
MAE and RMSE both measure errors relative to expert scores (lower is better) but RMSE penalizes large mistakes more. WTA≤1 measures the fraction of predictions that land within one point of the expert score. Bias measures the average signed error (systematic shift), with positive values indicating over-scoring and negative values indicating under-scoring. For ranking agreement within a problem, we use Kendall’s $\tau_{b}$ (ties-adjusted). Detailed definitions of Kendall’s $\tau_{b}$ as well as aggregation formulas are provided in §E.
@@ 3.3 What Factors Improve Single-pass Evaluator?
Evaluator MAE $\downarrow$ Bias $\approx$
o3 Gem. R1 o3 Gem. R1
[…]
Table 3: Per-generator evaluation accuracy. Red cells indicate highest MAE and highest positive bias for each evaluator.
@@
A natural question that can arise is the potential bias when evaluators assess outputs generated by their own models. To investigate this, we evaluate three single-pass evaluators, each using the same reference solutions and marking scheme, on per-generator splits, i.e., response sets grouped by their generators (R1, Gemini, o3). The results in Table 3 are mixed. On one hand, the strongest evaluator remains consistent across all generators: the o3-based evaluator. On the other hand, each evaluator exhibits its highest MAE on outputs produced by its own model family, indicating a tendency toward within-generator underperformance. Gemini and R1 exhibit a tendency to overscore their own responses, a bias not observed in the o3 evaluator.
@@ 3.4 Do Ensembling and Staged Workflows Improve Single-pass Evaluation?
We next extend single-pass evaluators through either ensembling or multi-stage pipelines.